Grade 11Physical Chemistry

an aques solution of sucrose has been labeled as 1 molal. find the mole fraction of solute and solvent

mateeullah
8 Years agoGrade 11

1 Answer

Ananya
8 Years ago
1 molal means 1 mole solute in 1000gm solvent
Moles of solvent=1000/18
=55.55
Mole fraction of solute =1/1+55.55
=0.177
Mole fraction of the solvent =1-0.177
=0.833